OpenStudy (anonymous):

'Equations on Graph Quiz' Explain your answer please The points on the graph are: (1/2,3) and (-4,-1) Which is an equation of the given line in standard form? A. –8x + 9y = 23 B. –8x + 9y = –23 C. –8x + 7y = 25 D. –9x + 8y = –23 Thank you for helping me, I'm new and I really appreciate it :)

imqwerty (imqwerty):

if u have any two points (x1,y1) and (x2,y2) then just use the point point form to get the eq of line- \[\frac{ y-y_{1} }{ y_{2}-y_{1} }=\frac{ x-x_{1} }{ x_{2}-x_{1} }\]

imqwerty (imqwerty):

*the 2 point form

OpenStudy (anonymous):

Thanks!

imqwerty (imqwerty):

np now u try to solve it and simplify